OpenStudy (anonymous):

a+b=10 a-b=2 Solve the system of equations

OpenStudy (anonymous):

add them together

OpenStudy (texaschic101):

a + b = 2 a - b = 2 -----------add 2a = 4 a = 4/2 a = 2 now sub 2 in for a in either of the above equations a + b = 2 2 + b = 2 b = 2 - 2 b = 0 check... a - b = 2 2 - 0 = 2 2 = 2 (correct) check... a + b = 2 2 + 0 = 2 2 = 2 (correct) ANSWER : a = 2 and b = 0